# Manual2Skill: Learning to Read Manuals and Acquire Robotic Skills for Furniture Assembly Using Vision-Language Models

### v1

@@ Abstract
Humans possess an extraordinary ability to understand and execute complex manipulation tasks by interpreting abstract instruction manuals. For robots, however, this capability remains a substantial challenge, as they cannot interpret abstract instructions and translate them into executable actions. In this paper, we present Manual2Skill, a novel framework that enables robots to perform complex assembly tasks guided by high-level manual instructions. Our approach leverages a Vision-Language Model (VLM) to extract structured information from instructional images and then uses this information to construct hierarchical assembly graphs. These graphs represent parts, subassemblies, and the relationships between them. To facilitate task execution, a pose estimation model predicts the relative 6D poses of components at each assembly step. At the same time, a motion planning module generates actionable sequences for real-world robotic implementation. We demonstrate the effectiveness of Manual2Skill by successfully assembling several real-world IKEA furniture items. This application highlights its ability to manage long-horizon manipulation tasks with both efficiency and precision, significantly enhancing the practicality of robot learning from instruction manuals. This work marks a step forward in advancing robotic systems capable of understanding and executing complex manipulation tasks in a manner akin to human capabilities.
@@ I Introduction
  • We propose Manual2Skill, a novel framework that leverages a VLM to learn complex robotic skills from manuals, enabling a generalizable assembly pipeline for IKEA furniture.
  • We introduce a hierarchical graph generation pipeline that utilizes a VLM to extract structured information for assembly tasks. Our pipeline facilitates real-world assembly and extends to other assembly applications.
@@
  • We perform extensive experiments to validate the effectiveness of our proposed system and modules.
@@ IV-A2 Stage I: Associating Real Parts with Manuals
Given the manual’s cover sketch of the assembled furniture and the pre-assembly scene image, the VLM aims to associate physical parts with the manual. The VLM achieves this by predicting the roles of each physical part through semantically interpreting the manual’s illustrations. This process involves analyzing spatial, contextual, and functional cues in the manual illustrations to enable a comprehensive understanding of each physical part. This design mimics human assembly cognition—people first map abstract manual images to physical parts before assembling. Our method follows CoT [51] and Least-to-Most [63] prompting, reducing cognitive load and improving accuracy. We considered pairwise matching of parts from manuals and scene images, but we found it impractical because the manuals not depict each part independently.
@@
For example, in Figure 2 In Stage I Output, we describe the chair’s seat as name: seat frame, label: [2], role: for people sitting on a chair, the seat offers essential support and comfort and is positioned centrally within the chair’s frame.. Here, [2] indicates that this triplet corresponds to the physical part labeled with index 2 in the pre-assembly scene image. This triplet format enhances interpretability and ensures consistency by structuring all outputs into the same data format. We use the Image Set and Text Instructions as the input prompt for the VLM (specifically GPT-4o [1]) and query it once to generate real assignments for all physical parts. We then use these labels as leaf nodes in the hierarchical assembly graph.
@@ IV-A3 Stage II: Identify Involved Parts in Each Step
In practice, we observe that irrelevant elements in the manual (e.g., nails, human figures) can distract the VLM. Following [49], we manually crop the illustrated parts and subassemblies in each manual step to focus the VLM’s attention (Figure 2 Stage II Image Set), significantly improving performance (see Ablation Study for details). Automating Region-of-Interest (ROI) detection remains an open problem beyond the scope of this work and is left for future research.
@@ IV-B2 Loss Function
We adopt a loss function that jointly considers pose prediction accuracy and point cloud alignment, following [60, 30]. The first term penalizes errors in the predicted SE(3) transformation, while the second measures the distance between predicted and ground truth point clouds. To account for interchangeable components, we compute the loss across all possible permutations of equivalent parts and select the minimum loss as the final training objective. We provide further details on the loss formulation and training strategy in  Section -B.
@@ V-A Hierarchical Assembly Graph Generation
TABLE I: Assembly Plan Generation Results.
Method Precision Recall F1 Score Success Rate
SingleStep 0.220 0.220 0.220 0.220
GeoCluster 0.197 0.201 0.196 0.080
Ours 0.690 0.680 0.684 0.620
@@
Experiment Setup. The IKEA-Manuals dataset [49] includes 102 furniture items, each with IKEA manuals, 3D parts, and assembly plans represented as trees in nested lists. We load each item’s 3D parts into Blender and render an image of the pre-assembly scene. Moreover, we split the 102 furniture items into two sets. The first set consists of 50 furniture items with six or fewer parts, and the second set contains 52 furniture items with seven or more parts. We observe that current VLMs can effectively deal with the first set, and a significant portion of real-world furniture also contains fewer than seven parts (as seen in real-world experiments). Here, we report the results of the first set. Please refer to Sections -D and -K for complete results and prompts. This rendered image, along with the manual, is processed by the VLM through the stages outlined in Section IV-A to generate a hierarchical assembly graph. Since we represent our graph as a nested list, we align our notation with the assembly tree notation used in IKEA-Manuals [49]. In this subsection, we refer to our generated assembly graph as the predicted tree.
Evaluation Metrics. We use the same metrics as IKEA-Manuals [49], which include precision, recall, and F1 score to compare predicted and ground-truth nodes of the assembly tree. For detailed descriptions of these metrics, we refer readers to [49].
The Matching criterion for each node is defined as follows: We consider a predicted non-leaf node correct only if its set of leaf and non-leaf child nodes exactly matches that of the corresponding ground-truth node(With consideration of equivalent parts). In other words, the predicted node must have the same children as its ground-truth counterpart. We compute precision, recall, and F1 scores based on this criterion.
The Success Rate criterion measures the proportion of the predicted tree that exactly matches the ground-truth tree. We consider a predicted tree exactly matched if all its non-leaf nodes satisfy the Matching criterion.
@@
Results. As shown in Table I, quantitative results demonstrate that both baseline methods face challenges in generating accurate assembly trees under the Matching and Assembly criterion. In contrast, our VLM-guided method achieves significantly superior performance, with a success rate of 62%. These findings underscore the robust generalization capabilities when guided by well-structured prompts. Figure 3 provides qualitative results for two furniture items, illustrating the advantages of our approach in greater detail. With the ongoing development of more advanced VLMs, we expect further enhancements in assembly planning accuracy. Please refer to Section -E for ablation results.
@@ V-B Per-step Assembly Pose Estimation
TABLE II: Qualitative Results of Pose Estimation.
@@ VI Limitations
This paper explores the acquisition of complex manipulation skills from manuals and introduces a method for automated IKEA furniture assembly. Despite this progress, several limitations remain. First, our approach mainly identifies the objects that need assembly but overlooks other details, such as grasping position markings and precise connector locations (e.g., screws). Integrating a vision-language model (VLM) module to extract this information could significantly enhance robotic insertion capabilities. Second, the method does not cover the automated execution of fastening mechanisms, like screwing or insertion actions, which depend heavily on force and tactile sensing signals. We leave these challenges as directions for future work.

### v2

@@ Abstract
Humans possess an extraordinary ability to understand and execute complex manipulation tasks by interpreting abstract instruction manuals. For robots, however, this capability remains a substantial challenge, as they cannot interpret abstract instructions and translate them into executable actions. In this paper, we present Manual2Skill, a novel framework that enables robots to perform complex assembly tasks guided by high-level manual instructions. Our approach leverages a Vision-Language Model (VLM) to extract structured information from instructional images and then uses this information to construct hierarchical assembly graphs. These graphs represent parts, subassemblies, and the relationships between them. To facilitate task execution, a pose estimation model predicts the relative 6D poses of components at each assembly step. At the same time, a motion planning module generates actionable sequences for real-world robotic implementation. We demonstrate the effectiveness of Manual2Skill by successfully assembling several real-world IKEA furniture items. This application highlights its ability to manage long-horizon manipulation tasks with both efficiency and precision, significantly enhancing the practicality of robot learning from instruction manuals. This work marks a step forward in advancing robotic systems capable of understanding and executing complex manipulation tasks in a manner akin to human capabilities. Project Page: https://owensun2004.github.io/Furniture-Assembly-Web/
@@ I Introduction
  • We propose Manual2Skill, a novel framework that leverages VLM to learn robotic skills from manuals, enabling a generalizable assembly pipeline for IKEA furniture.
  • We introduce a hierarchical graph generation pipeline that utilizes VLM to extract structured information for assembly tasks. Our pipeline facilitates real-world assembly and extends to other assembly applications.
@@ IV-A2 Stage I: Associating Real Parts with Manuals
Given the manual’s cover sketch of the assembled furniture and the pre-assembly scene image, the VLM aims to associate physical parts with the manual. The VLM achieves this by predicting the roles of each physical part through semantically interpreting the manual’s illustrations. This process involves analyzing spatial, contextual, and functional cues in the manual illustrations to enable a comprehensive understanding of each physical part. Our method follows CoT [51] and Least-to-Most [63] prompting for better accuracy.
@@
For example, in Figure 2 In Stage I Output, we describe the chair’s seat as name: seat frame, label: [2], role: for people sitting on a chair, the seat offers essential support and comfort and is positioned centrally within the chair’s frame. Here, ‘[2]’ indicates that this triplet corresponds to the physical part labeled with index 2 in the pre-assembly scene image. This triplet format enhances interpretability and ensures consistency by structuring all outputs into the same data format. We use the Image Set and Text Instructions as the input prompt for the VLM (specifically GPT-4o [1]) and query it once to generate real assignments for all physical parts. We then use these labels as leaf nodes in the hierarchical assembly graph.
@@ IV-A3 Stage II: Identify Involved Parts in Each Step
In practice, we observe that irrelevant elements in the manual (e.g., nails, human figures) can distract the VLM. We use the cropped manual images from [49], where only the furniture parts and subassemblies are reserved to focus the VLM’s attention (Figure 2 Stage II Image Set), significantly improving performance (see Section -E for details).
@@ IV-B2 Loss Function
t jointly considers pose prediction accuracy and point cloud alignment, following [60, 30]. The first term penalizes errors in the predicted SE(3) transformation, while the second measures the distance between predicted and ground truth point clouds. To account for interchangeable components, we compute the loss across all possible permutations of equivalent parts and select the minimum loss as the final training objective. We provide further details on the loss formulation and training strategy in  Section -B.
@@ V-A Hierarchical Assembly Graph Generation
TABLE I: Success Rate Across Task Complexity ($\uparrow$)
Number of Parts 2$\sim$4 5$\sim$6 7$\sim$8 9+ Average
SingleStep 35.7 16.7 0 0 10.8
GeoCluster 21.4 2.8 0 0 3.9
Ours (Scene Variations) 78.6 50.0 12.5 0 31.4
Ours 78.6 55.6 8.3 0 32.4
Furniture Count 14 36 24 28
@@
Figure 4: Pre-Assembly Scene Variations. (Left) original pre-assembly scene. (Middle) parts randomly shuffled along the ground plane. (Right) parts randomly rotated in-place.
Experiment Setup. The IKEA-Manuals dataset [49] includes 102 furniture items, each with IKEA manuals, 3D parts, and assembly plans represented as trees in nested lists. For each item, we load its 3D parts into Blender and render two images: one depicting the original pre-assembly scene with neatly arranged parts, and another showing a scene variation where parts are arbitrarily perturbed (e.g., rotated and shuffled) along a ground plane (see Figure 4). This randomization introduces diversity, better simulating real-world pre-assembly scenarios where parts may be disorganized.
Each rendered image of the pre-assembly scene, along with the manual, is processed by the VLM through the stages outlined in Section IV-A to generate a hierarchical assembly graph. Since we represent our graph as a nested list, we align our notation with the assembly tree notation used in IKEA-Manuals [49]. In this subsection, we refer to our generated assembly graph as the predicted tree.
Evaluation Metric. We use Success Rate criterion, which measures the proportion of the predicted tree that exactly matches the ground-truth tree. We consider a predicted tree exactly matched if all its nodes contain the same set of children nodes as their ground-truth counterparts. All equivalent parts are seen as identical when computing the tree matching.
@@
Results. As shown in LABEL:tab:assembly_plan, understanding and interpreting manuals is a challenging task. Our VLM‑guided approach effectively handles furniture manuals with up to 6 parts, a significant breakthrough where baselines struggle even in simpler cases. This $\leq 6$ parts threshold reflects current VLM capabilities in complex visual-spatial reasoning tasks. Our framework is designed to be scalable, benefiting from rapid VLM advancements. We anticipate performance improvements as more powerful VLMs emerge. LABEL:tab:assembly_plan also highlights the VLM’s generalization ability: pre-assembly scene variations in Figure 4 minimally impact assembly graph generation performance, with average success rates drop within 1% across all 102 furniture items. Figure 3 provides qualitative results for two furniture items, illustrating the advantages of our approach in greater detail.
For ablation studies on the necessity of using and segmenting manuals, see Section -E. Failure cases are further explained in Section -F. Additional results and prompt templates are in Sections -D and -K, respectively.
@@ V-B Per-step Assembly Pose Estimation
TABLE II: Quantitative Results of Pose Estimation.
@@ VI Limitations
This paper explores the acquisition of complex manipulation skills from manuals and introduces a method for automated IKEA furniture assembly. Despite this progress, several limitations remain. First, our approach mainly identifies the objects that need assembly but overlooks other details on the manual, such as grasping position markings and precise connector locations (e.g., screws). Integrating a vision-language model (VLM) module to extract this information could significantly enhance robotic insertion capabilities. Second, the method does not cover the automated execution of fastening mechanisms, like screwing or insertion actions, which depend heavily on force and tactile sensing signals. We leave these challenges as directions for future work.